OpenStudy (anonymous):

Verify the identity cos 4u= cos^2u - sin^2u

OpenStudy (luigi0210):

You'll have to use the identities \(\large sin^2x= \frac{1}{2}(1-cos2x) \) and \(\large cos^2 x = \frac{1}{2}(1+cos2x)\)

OpenStudy (anonymous):

Sorry it's Cos 4u= cos^2(2u) -sin^2 (2u)

OpenStudy (anonymous):

This is the double angle formula

OpenStudy (anonymous):

multiplied by a factor of 2

OpenStudy (michele_laino):

Please keep in mind this identity: \[\cos \left( {2x} \right) = {\left( {\cos x} \right)^2} - {\left( {\sin x} \right)^2}\]

OpenStudy (anonymous):

OK so do I multiply everything by 2?

OpenStudy (anonymous):

Would cos2 (2×) = 2 (cos x)^2 - 2 (sin x)^2 be right?

OpenStudy (michele_laino):

Please set x= 2u into my above identity

OpenStudy (anonymous):

cos2 (2u) = (cos 2u)^2 - (sin 2u)^2

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

yes! it is your dentity, since 2*2u=4u

OpenStudy (michele_laino):

oops..identity...

OpenStudy (anonymous):

Oh so that's the answer? There's nothing else I need to do?

OpenStudy (michele_laino):

yes! that is the answer since you have showed that your originalk identity is true, and in order to that you ahve applied the subsequent identity: \[\cos \left( {2x} \right) = {\left( {\cos x} \right)^2} - {\left( {\sin x} \right)^2}\]

OpenStudy (michele_laino):

oops...you have...

OpenStudy (anonymous):

OK thank you!

OpenStudy (michele_laino):

thank you!